OpenStudy (anonymous):

my question got answered, bu could some explain to me how they arrived at the asnwer. I qwas able to get the first part, 1.1, then i mutpolied it by 4 and got something way differet.

OpenStudy (anonymous):

ok well u said that 4 was an exponent - an exponent is a power or it somteimes caled an index for example 3^3 is '3 to the power 3' 3 is the exponent it means ' 3 times 3 times 3' = 27. In same way 4^2 = 4 x 4 = 16. 6^4 = 6 x 6 x 6 x 6 . so 1.1 ^ 4 = 1.1 x 1.1 x 1.1 x 1.1

OpenStudy (anonymous):

1.1 ^ 4 = 1.4641

OpenStudy (anonymous):

ok, I get it. But is here an easier way to figure that out. Is there something on the calculator that will do this? Why is it I dont get the same answer when I multiplied 1.1 by 4?

OpenStudy (anonymous):

On my calculator I go 1.1 SHIFT (x^y) 4 (1.1 is x, 4 is y) to get 1.4641. 1.1 multiplied by 4 = 1.1 + 1.1 + 1.1 + 1.1 1.1 to power 4 = 1.1 * 1.1 *1.1 * 1.1

OpenStudy (anonymous):

Thank you sooooo much. You are awesome. I have been racking my brain to figure this out.